UT 20171029 LUCI1/2 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt (LSW), DThompson (LBTO)
Observer AO-Team: DMiller, GTaylor (LBTO)
Telescope Operator: DHuerta

Observations done remotely from Tucson

Plan for the night

- LUCI1/2 AFC AO-ima commissioning in binocular mode

- If bino mode not successfull, revert to monocular mode, 1st priority for LUCI2

- If seeing is bad, handover to Olga for MODS checkout

Summary

- Excellent night, AO-weather all night

- AO-com AFC-imaging at low elevation completed

- AOs worked perfectly for almost 5h in parallel during 2nd part of night

- Lost about 3.5hrs due to frequent errors on both AOs during 1st part of night

Details

All times are UT
Files are luci1/2.20171029.NNNN.fits

Doug: AO-man @ SX
Greg: AO-man @ DX

00:37 Sunset

Weather: Clear, T = +7deg, hum = 50%, very low wind

00:55 David opens up, checks pointing and collimates

At the same time SHooper kicked off a couple of people logged in at the obs machines,
as the desktop onto the mountain here was extremely slow

01:35 Doug & Greg make sanity check of the AO-systems as ARGOS had lots of troubles
last night with their system

Bayside stage @ SX had to be homed (manually at rotangle 341)

DIMM reports below 1"

No NCPA @ SX, no signal for the optical gain there

AFC-imaging, binocular mode

02:19 Preset to M15, OK, RunAO, skip frame => caused shell to RIP @ SX, loop closed @ DX,
reason: 3 stars nearby (we are in a globular cluster) so light from them was spilling on the WFS,
Doug tries to close the loop manually, did it, but opened quickly afterwards

02:52 Resend preset again, OK, RunAO, ICE communication error @ DX

02:56 Resend preset again, OK, RunAO, finally on both sides closed @ 03:04,
sent offset 0,0, AO@SX start to behave supercrappy.., Doug is fighting with it

03:29 David is correcting the pointing (that can be the reason for the issues with the AO@SX)

03:38 Resend preset again, OK, RunAO, helped in principle but issue returned at SX

We give up, AO@SX is not feasible in this field, AO@DX does not have that problem
for more information, check Dougs AO-logs

03:49 Preset to M2, OK, RunAO, OK, AFC-alignment, OK @ LUCI2, bad @ LUCI1
positions of holes must newly be defined and included, will continue to run in bino

GCS grabbed different stars for guiding and collimation which lead to two
different AO-ref stars (we were lucky to be in a globular cluster)
SX is at bin=3, DX is at bin=2

=> we need to check and correct pointing before every preset !

AFC@DX looks great (ima's #9-11, though the wings look slightly elongated)

On1min comparison images without AFC but with passive flexure compensation (ima's #12-16)
the stars are strongly elongated at PA = -45deg on both sides. Called JHill to see
whats the reason is.

Continue with AFC H-band: #19-20, and comparison images 1min without any compensation (#23)
and passive flexure compensation (#24). Elongation has disappeared. According to JHill no
strong vibrations are present.

05:25 Preset to Mayall 2, OK, RunAO, OK, SRR alignment OK, start SRR, error AO@DX (human error),
known error AO@SX (stage moved because of offset but telescope not, offset 0,0)

05:59 Resent preset to Mayall 2 again, OK, RunAO, OK (bin=4, 200 Hz), SRR alignment OK,

Ks: #30-32 (SRR), #33-37 (short), ele: 79 - 74
H: #39-41 (SRR), #41 looks bad, #42-43 (short), offset error @DX after #43,
took Greg quite a while to sort that out, reason was a communication error, WUnit etc. not responsive

07:24 New preset to Mayall 2 required, skip the remaining H-band short images,
preset failed on DX, AdSec had to be restartet, and prim.&sec. optics be cleared

07:44 Preset once more to Mayall 2, OK, RunAO OK, start SRR, OK

J: #47-49 (SRR), OK;, #50 (short), than offset failed because loop opened at the end
of #50 (not due to seeing),
, recovered, #51-54 OK, ele: 57

DIMM ~ 0.8"

Br_gam: #57-59 (SRR), OK smile 60-64 (short)

Fe II: #67-69 (SRR), OK, smile #70-74 (short)

P_beta: #77-79 (SRR), OK, smile #80-84 (short)

DIMM ~ 0.9"

OK; 1st pass in all filters with LUCI2 at high elevation (K), medium (HJ), and low (Br_gam, Fe II, P_beta) successfull

Time to complete about 4h 40min (120min on target)

Alexander has implemented new AFC-hole positions for LUCI1, do full bino-AFC now

10:10 Preset to Theta Orionis C, OK, RunAO, OK, align AFC, OK

Ks:
L1: #37-39 (SRR) OK, #40-44 (short)
L2: #89-91 (SRR) OK, # 92-96 (short)

H:
L1: #47-49 (SRR), #50-54 (short)
L2: #99-101 (SRR), #102-106 (short)

J: dropped to allow completion of narrow-band filters

Br_gam:
L1: #58-60 (SRR), #61-65 (short)
L2: #109-111 (SRR), #112-116 (short)

DIMM 0.6"

Fe II:
L1: #67-69 (SRR), #70-74 (short)
L2: #118-120 (SRR), #121-125 (short)

P_beta:
L1: #79-81 (SRR), 82-86 (short)
L2: #130-132 (SRR), 133-137 (short), #130 slight AFC-shift

Time to completion for 5 filters (3 * 5min SRR + 5 * 1min per filter): 2h 35m (1h 40m on target)

OK; pass in all filters with LUCI1+LUCI2 at low elevation in K H Br_gam Fe II und P_beta successfull

WOW, both AOs worked almost 5h without any issues in parallel, super !!!!!

12:44 Offset failed @ SX (after ima #84) at the very last image of the night, bad luck frown, sad smile

12:28 End of observing

Great night

-- %USERSIG{JochenHeidt - 2017-10 -29}%
Topic revision: r7 - 31 Oct 2017, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback